# Deep neural networks for choice analysis: Enhancing behavioral regularity with gradient regularization

### v1

@@ Abstract
Deep neural networks (DNNs) have been increasingly applied to travel demand modeling because of their automatic feature learning, high predictive performance, and economic interpretability. However, DNNs frequently present behaviorally irregular patterns, significantly limiting their practical potentials and theoretical validity in travel behavior modeling. This study proposes strong and weak behavioral regularities as novel metrics to evaluate the monotonicity of individual demand functions (a.k.a. “law of demand”), and further designs a constrained optimization framework with six gradient regularizers to enhance DNNs’ behavioral regularity. The empirical benefits of this framework are illustrated by applying these regularizers to travel survey data from Chicago and London, which enables us to examine the trade-off between predictive power and behavioral regularity for large versus small sample scenarios and in-domain versus out-of-domain generalizations. The results demonstrate that, unlike models with strong behavioral foundations such as the multinomial logit, the benchmark DNNs cannot guarantee behavioral regularity. However, after applying gradient regularization, we increase DNNs’ behavioral regularity by around 6 percentage points while retaining their relatively high predictive power. In the small sample scenario, gradient regularization is more effective than in the large sample scenario, simultaneously improving behavioral regularity by about 20 percentage points and log-likelihood by around 1.7%. Comparing with the in-domain generalization of DNNs, gradient regularization works more effectively in out-of-domain generalization: it drastically improves the behavioral regularity of poorly performing benchmark DNNs by around 65 percentage points, indicating the criticality of behavioral regularization for enhancing model transferability and application in forecasting. Moreover, the proposed optimization framework is applicable to other neural network–based choice models such as TasteNets. Future studies could use behavioral regularity as a metric along with log-likelihood, prediction accuracy, and $F_{1}$ score in evaluating travel demand models, and investigate other methods to further enhance behavioral regularity when adopting complex machine learning models.
Keywords: travel demand, deep learning, behavioral regularization
@@ Introduction
Deep neural networks (DNNs) have revolutionized fields such as computer vision and natural language processing, that in turn support technologies such as self-driving cars and large language models (van Dis et al., 2023; LeCun et al., 2015). DNNs have also been applied to economics (Zheng et al., 2023), including interpreting and predicting individual choice behavior (Wang et al., 2020a, b). It is in this latter area that DNNs offer a contrast with the conventionally used discrete choice models (DCMs), which are typically based on random utility maximization (Ben-Akiva and Lerman, 1985) and the assumption that travelers choose the alternative with the highest random utility in the choice set. One drawback of this traditional modeling paradigm is the time-consuming trial-and-error process for an “optimal” specification of the model, in particular the utility function (van Cranenburgh et al., 2022), which represents economic preferences. Additionally, the decisions made in this process are often subjective. By contrast, DNNs are capable of automated feature learning, i.e., the specification of a DNN-based choice model is automatically learned from the input data, which avoids the aforementioned specification search process, and reduces the level of subjectivity. The high prediction accuracy of DNNs is a result of their complex model structure, which helps capture intricate behavioral relationships and can provide new insights beyond those of conventional DCMs.
@@
Traditional choice modelers often see DNNs as “black-box” models, although DNNs actually contain complete economic information for choice analysis (Wang et al., 2020b). However, existing DNNs often exhibit behaviorally irregular patterns because the demand functions in DNNs are not guaranteed to decrease monotonically with generalized costs. The “law of demand” in economics indicates an inverse relationship between generalized costs and the aggregate demand. While DCMs such as random utility models (RUMs), do not impose specific directionalities a priori, the specification search conducted by an analyst will not accept models that lead to counter-intuitive results. With DNNs, the analyst has less control, and non-monotonic patterns have been detected empirically in DNNs’ predictions, even with model ensembles (Wang et al., 2020a, b; Xia et al., 2023). This fact might be a drawback of the nonlinear structure of DNNs, making them flexible to fit data but difficult to restrict the gradient’s direction with limited data samples. The issue often deteriorates in case of out of sample application, i.e., applying a trained DNN to a testing set with unseen distributions (Quiñonero-Candela et al., 2008). In fact, the out-of-domain generalizability of DNNs has attracted rising interests in several computer science fields, including domain adaption (Wang and Deng, 2018) and transfer learning (Pan and Yang, 2009).
@@
To improve the monotonicity of DNNs, we propose to regularize the loss function in training, which has been shown in computer science to enhance the robustness of DNNs (Lyu et al., 2015; Ross and Doshi-Velez, 2018). However, this approach has rarely been considered in previous DNN-based choice models, or used for enhancing behavioral regularity (Wang et al., 2020b; Zheng et al., 2021). In this paper, we address the issue of behavioral irregularity by first defining strong and weak behavioral regularity metrics based on monotonicity of the demand functions, and further designing and implementing a constrained optimization framework that regularizes the input gradient in order to explicitly constrain the gradient’s direction. We then design experiments to examine the performance of behaviorally regularized DNNs in terms of behavioral regularity and predictive performance, differentiating the in-domain versus out-of-domain generalization. We also consider factors such as sample size, which is valuable for practical modeling because large samples are costly for travel surveys. Based on two travel survey datasets from Chicago and London, our experiments compare across the standard DNN and TasteNet (Han et al., 2022) architectures using five evaluation metrics, including log-likelihood, prediction accuracy, $F_{1}$ score, strong behavioral regularity, and weak behavioral regularity. The multinomial logit (MNL) is chosen as a benchmark model due to its concise expression and high behavioral regularity. The results show that by using appropriate gradient regularization, both DNNs and TasteNets can achieve high regularity without sacrificing their predictive power, which makes these models competitive in real-world applications and demonstrates the generality of our gradient regularization framework.
@@
The economic choice behavior of humans generally follows the “law of demand” in economics, which states the inverse relationship between price and quantity demanded. This law leads to a monotonic change in market demand due to the change in consumers’ purchasing power, including price and income changes (Chiappori, 1985; Härdle et al., 1991; Hildenbrand, 1983; Quah, 2000). The transportation field has also observed the negative influence of travel costs on travel demand (McFadden, 1974; Souche, 2010; Yao and Morikawa, 2005), based on which demand management policies such as road pricing (May, 1992; Yang and Bell, 1997) were developed. Although such market rationality is widely recognized, individual choice behavior might be irrational (Becker, 1962; Knez et al., 1985). Lichtenstein and Slovic (1971) studied preference reversal in decision making, which is a typical counterexample of individual rationality. Studies in bounded rationality theory (Simon, 1957; Di and Liu, 2016; Watling et al., 2018) and prospect theory (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992) also relax the strict monotonicity assumption in modeling demand.
@@
The aforementioned law is generally followed by the design of random utility models. In the MNL model, for example, an increase in the travel cost of an alternative would be expected to decrease its systematic utility, thus decreasing its choice probability by design. If initial model estimation leads to counter-intuitively signed coefficients, such as positive cost coefficients, then this is easily spotted by an analyst and serves as an invitation to refine the model specification or deal with data issues. Once all signs are as expected, the monotonic relationship is guaranteed. By contrast, this is not the case in DNNs because of the complex nonlinear model structure, especially when the number of hidden layers increases. For example, Xia et al. (2023) observed non-monotonic demand predictions with increasing generalized costs in a mode choice experiment with DNNs, which suggests the need to investigate the monotonicity of DNNs to improve their behavioral regularity. Although shallow neural networks (NNs) might reduce the risk of non-monotonic behavior of DNNs (Alwosheel et al., 2019; Zhao et al., 2020), this might come at the cost of reduced modeling flexibility and universal approximation power. Alternatively, Han et al. (2022) and Sifringer et al. (2020) proposed to use DNNs only for learning latent representation in the utility function, while resorting to the DCM framework to ensure model monotonicity. For example, TasteNet, the neural-embedded DCM proposed by Han et al. (2022), assumes linear model specification, while coefficients are parameterized by neural network. In this paper, the TasteNet is considered as a reference to the standard DNN architecture. On the other hand, depending on model design, these hybrid DNN models might still produce irregular predictions (Wang et al., 2021; Wong and Farooq, 2021). Moreover, hybrid DNN models are a compromise for regularity since they again require the subjective process of model specification. To fully utilize the capability of DNNs, previous studies have attempted to migrate the non-monotonic issue through model ensemble. However, irregular patterns might still be observed after averaging over multiple trainings (Wang et al., 2020a, b; Xia et al., 2023). One promising direction is to directly integrate domain-specific knowledge into the design and training of DNNs, such as incorporating monotonicity constraints in model training (Haj-Yahia et al., 2023). However, there is no consensus on how to measure or improve the model regularity of DNNs within the choice modeling field. This paper contributes to the development of a behavioral regularity measure and a novel regularization framework.
@@
As discussed in computer science applications, the regularity of DNNs can be improved by employing either hard or soft constraints. The first category enforces monotonicity by model construction, e.g., constraining the positiveness of weights in hidden layers (Daniels and Velikova, 2010; Dugas et al., 2009; Sill, 1997) through non-negativity constraint (Lawson and Hanson, 1995), restricting the derivative to be positive (Neumann et al., 2013; Wehenkel and Louppe, 2019), down-weighting samples that violate monotonicity (Archer and Wang, 1993), and incorporating deep lattice network for learning monotonic functions (You et al., 2017). The second category achieves monotonicity by regularization, i.e., by augmenting a regularization term in the loss function to jointly improve model monotonicity. Regularization is firmly rooted in constrained optimization, including the Lagrangian method as an example (Boyd and Vandenberghe, 2004). It has been widely applied as a local method to penalize constraint violations. For example, Sill and Abu-Mostafa (1996) penalized squared deviations in monotonicity for virtual pairs of input variables, while Gupta et al. (2019) proposed a pointwise loss that embeds prior knowledge about monotonicity. Moreover, gradient regularization has also been used to enhance model robustness against adversarial examples (Lyu et al., 2015; Ross and Doshi-Velez, 2018), e.g., penalizing the squared $L_{2}$ norm of the gradient of the loss with respect to (w.r.t.) inputs (Drucker and Le Cun, 1991; Ororbia II et al., 2017), penalizing the squared Frobenius norm of the Jacobian matrix of probabilities (Sokolić et al., 2017) and utilities (Jakubovitz and Giryes, 2018) w.r.t. inputs. Note that regularizing the gradient norms might be less effective to improve monotonicity than regularizing the gradient’s direction like Haj-Yahia et al. (2023). Inspired by the aforementioned regularization methods, we will design our own approaches in the demand modeling context.
@@
The discrete choice problem is cast as a supervised classification problem in DNN-based choice analysis. Assuming there are in total $D$ explanatory variables $(x_{1},\ldots,x_{D})$ for all alternatives, the attribute vector of individual $n$ can be written as $\mathbf{x}_{n}=[x_{n1},\ldots,x_{nD}]^{\top}$. Then, a DNN model predicts the probability of $n$ choosing $i$ out of $J$ alternatives, i.e., $P_{ni}:\mathbb{R}^{D}\to(0,1)$ and $\sum_{i=1}^{J}P_{ni}=1$. The observed choice vector $\mathbf{y}_{n}\in\{0,1\}^{J}$ of $n$ is used for DNN training, where $y_{ni}=1$ if alternative $i$ is chosen, and $y_{ni}=0$ otherwise. Similar to conventional RUMs, DNN models aim to find specifications with high predictive power and behavioral regularity.
However, contrasting to the manual model specifications of RUMs, DNNs automatically learn model specifications with their unique representation learning capability. Specifically, utility vector $\mathbf{V}_{n}=[V_{n1},\ldots,V_{nJ}]^{\top}$ is specified through a series of transformations, termed as layers $(f_{1},\ldots,f_{H})$, where $H$ denote the total number of layers in a DNN. Each layer $f_{h}$ contains a learnable parameter matrix $W_{h}$, a bias vector $\mathbf{b}_{h}$, and an activation function $\varphi(\cdot)$ (e.g., the rectified linear unit, ReLU) to transform $\mathbf{x}_{n}$. Specifically, each layer transformation can be written as
@@
where $W^{J\times D}$ can be interpreted as coefficients and $\mathbf{b}^{J\times 1}$ as alternative-specific constants. Although closely related to the MNL model, DNNs allow flexible model specification through multi-layer nonlinear transformations. We illustrate in Fig. 1 a feedforward DNN structure with four hidden layers and one classification layer for a choice modeling problem with $D$ attributes for $J$ alternatives.
@@
In this study, we propose a novel metric to evaluate behavioral regularity, which measures the monotonicity of the aggregate choice probability functions. The proposed metric essentially measures the monotonicity consistency between the model and prior knowledge on the correct signs of parameter estimates, commonly used in the subjective process for selecting plausible specification of RUMs. We define the behavioral regularity metric of alternative $i$ w.r.t. a cost variable $x_{d}$ as
@@
where $z$ represent the characteristic factor of a population group, $P_{i}(\mathbf{x}_{z})$ maps the individual’s attributes $\mathbf{x}_{z}$, including both individual-specific sociodemographic variables mapped from factor $z$, and alternative-specific cost variables, to the individual’s probability of choosing $i$, and $\mathbbm{1}\{\cdot\}$ is an indicator function that equals 1 if ${\partial P_{i}(\mathbf{x}_{z})}/{\partial x_{d}}<\varepsilon$, and 0 otherwise. The $\rho(x_{d},z)$ term denotes the joint density of attributes $x_{d}$ and the population factor $z$. Parameter $\varepsilon$ represents the modeler’s prior assumptions on the monotonicity of $P_{i}$ w.r.t. $x_{d}$:
  1. (1)
@@
  2. (2)
    $\varepsilon>0$, termed as weak regularity, relaxes the strict monotonicity assumption and allows $P_{i}$ to be non-decreasing w.r.t. $x_{d}$. The formulation assumes that some population groups $\{z\}$ do not respond (with zero derivative) to $x_{d}$, implying that the behavioral regularity constraint becomes weaker.
As an illustration, a classical linear MNL model with a negative parameter w.r.t. $x_{\cdot d}$ yields $B_{id}=1$, which implies that all individual behaviors are consistent with the demand monotonicity assumption and well-captured by the model. We also note that Eq. 5 can be interpreted as the cumulative distribution of behavioral regularity with $\varepsilon$ over the whole population and the domain of $x_{d}$.
The population-based behavioral regularity measure in Eq. 5 can be approximated by the mean behavioral regularity across individuals:
@@
where $N$ is the sample size, and the partial derivative is computed with finite differences. The proposed empirical regularity measure in Eq. 6 is the sample analog of the exact metric in Eq. 5. By the Glivenko–Cantelli theorem, Eq. 6 converges in probability to Eq. 5, the cumulative distribution function for the mean population behavioral regularity with $\varepsilon$.
Our behavioral regularity metrics can be extended to incorporate taste heterogeneity across population groups and even individuals within each group by distinguishing $\varepsilon$ w.r.t. different groups, that is, $\varepsilon$ can be further specified as $\varepsilon_{z}$ to reflect the group-specific thresholds. Meanwhile, our behavioral regularity metrics $B_{id}$ only require the aggregate regularity rather than individual one, which is inspired by classical economics discussions that market rationality is a fundamental law while individual behaviors might present more diverse and irrational patterns (Becker, 1962).
@@
The unconstrained formulation in Eq. 7 is sufficient for the estimation of conventional DCMs, since they often satisfy convexity conditions under linear utility specification. In a linear MNL model, for example, choice probability $P_{ni}$ increases monotonically with utility $V_{ni}$ according to Eq. 3. The linear specification in Eq. 4 induces monotonicity of utility w.r.t. cost variables. Therefore, if individuals indeed perceive higher utility with lower costs, the optimization would result in negative parameter estimates, and the behavioral monotonicity is clearly satisfied by evoking the chain rule. The multi-layer nonlinear transformations in NNs allow for approximation of arbitrary functions, but these complex transformations might lead to non-monotonic choice probability functions, especially when the network is deep. In this case, unconstrained likelihood maximization can no longer guarantee a behaviorally regular model.
@@
where $R$ constrains the attributes $\mathbf{x}_{n}$ of an individual $n$, defined as $R:\mathbb{R}^{J\times D}\to\mathbb{R}$, where dimension $J$ is the number of alternatives and $D$ is the number of attributes. Hence the behavioral regularity constraints are imposed at the individual level to achieve the aggregate behavioral regularity in $B_{id}$. The specific behavioral regularity constraints will be designed in the next subsection.
@@
Compared to the hard constraint formulation, soft regularization can flexibly accommodate the various degrees of validity in our behavioral regularity assumptions. Similar to the motivation for the weak regularity metric, our approach allows each individual $n$ to somewhat violate the preset constraint $R(\mathbf{x}_{n})$, thus accommodating the potentially irregular behavior of certain individuals. As a result, hyperparameter $\lambda$ provides insight into the consistency between behavioral regularity assumptions and the actual behavior of studied individuals. If a larger $\lambda$ is required to achieve higher predictive performance, it might imply that the actual behavior is inconsistent with prior assumptions, thus providing extra insight into the validity of behavioral regularity constraints.
@@
For individual $n$, the Jacobian matrix (gradient) of demand vector $\mathbf{P}=[P_{1},\ldots,P_{J}]^{\top}$ w.r.t. cost variables $\{x_{1},\ldots,x_{D}\}$ can be written as
@@
  1. (1)
    Direct derivatives: e.g., the probability of driving w.r.t. driving cost.
  2. (2)
    Cross derivatives: e.g., the probability of driving w.r.t. train time.
  3. (3)
    Sociodemographic derivatives: e.g., the probability of driving w.r.t. the age of traveler.

### v2

@@ Abstract
Deep neural networks (DNNs) have been increasingly applied in travel demand modeling because of their automatic feature learning, high predictive performance, and economic interpretability. Nevertheless, DNNs frequently present behaviorally irregular patterns, significantly limiting their practical potentials and theoretical validity in travel behavior modeling. This study proposes strong and weak behavioral regularities as novel metrics to evaluate the monotonicity of individual demand functions (known as the “law of demand”), and further designs a constrained optimization framework with six gradient regularizers to enhance DNNs’ behavioral regularity. The empirical benefits of this framework are illustrated by applying these regularizers to travel survey data from Chicago and London, which enables us to examine the trade-off between predictive power and behavioral regularity for large versus small sample scenarios and in-domain versus out-of-domain generalizations. The results demonstrate that, unlike models with strong behavioral foundations such as the multinomial logit, the benchmark DNNs cannot guarantee behavioral regularity. However, after applying gradient regularization, we increase DNNs’ behavioral regularity by around 6 percentage points while retaining their relatively high predictive power. In the small sample scenario, gradient regularization is more effective than in the large sample scenario, simultaneously improving behavioral regularity by about 20 percentage points and log-likelihood by around 1.7%. Compared with the in-domain generalization of DNNs, gradient regularization works more effectively in out-of-domain generalization: it drastically improves the behavioral regularity of poorly performing benchmark DNNs by around 65 percentage points, highlighting the criticality of behavioral regularization for improving model transferability and applications in forecasting. Moreover, the proposed optimization framework is applicable to other neural network–based choice models such as TasteNets. Future studies could use behavioral regularity as a metric along with log-likelihood, prediction accuracy, and $F_{1}$ score when evaluating travel demand models, and investigate other methods to further enhance behavioral regularity when adopting complex machine learning models.
Keywords: travel demand, deep learning, choice analysis, behavioral regularization
@@ Introduction
Deep neural networks (DNNs) have revolutionized fields such as computer vision and natural language processing, which in turn support technologies such as self-driving cars and large language models (van Dis et al., 2023; LeCun et al., 2015). DNNs have also been applied in economics (Zheng et al., 2023) to interpret and predict individual choice behavior (Wang et al., 2020a, b). It is in this latter area that DNNs offer a contrast with the conventionally used discrete choice models (DCMs), which are typically based on random utility maximization and the assumption that travelers choose the alternative with the highest random utility in the choice set (Ben-Akiva and Lerman, 1985). One drawback of this traditional modeling paradigm is the time-consuming trial-and-error process for an “optimal” specification of the model, in particular the utility function that represents economic preferences (van Cranenburgh et al., 2022). Additionally, the decisions made in this process are often subjective. By contrast, DNNs are capable of automatic feature learning, i.e., the specification of a DNN-based choice model is automatically learned from the input data, which avoids the specification search process and reduces the level of subjectivity. The high prediction accuracy of DNNs is a result of their complex model structure, which can capture intricate behavioral relationships and provide new insights beyond those of conventional DCMs.
@@
Traditional choice modelers often see DNNs as “black-box” models, although DNNs actually contain complete economic information for choice analysis (Wang et al., 2020b). However, existing DNNs often exhibit behaviorally irregular patterns because the demand functions in DNNs are not guaranteed to decrease monotonically with generalized costs. The “law of demand” in economics indicates an inverse relationship between generalized costs and the aggregate demand. While DCMs such as random utility models (RUMs) do not impose specific prior directionality, the specification search conducted by an analyst would prevent models that lead to counter-intuitive results. However, non-monotonic patterns have been detected empirically in DNNs’ predictions, even with model ensembles (Wang et al., 2020a, b; Xia et al., 2023), as the analyst has less control. This fact might be a drawback of the nonlinear structure of DNNs, allowing them to flexibly fit the data but making it challenging to constrain the gradient’s direction with limited data samples. The issue often deteriorates in out-of-domain applications, i.e., applying a trained DNN to a test set with unseen distributions (Quiñonero-Candela et al., 2008). In fact, the out-of-domain generalizability of DNNs has attracted rising interests in several computer science subfields, including domain adaption (Wang and Deng, 2018) and transfer learning (Pan and Yang, 2009).
@@
To improve the monotonicity of DNNs, we propose to regularize the loss function in training, which has been shown in computer science to enhance the robustness of DNNs (Lyu et al., 2015; Ross and Doshi-Velez, 2018). Nevertheless, this approach has rarely been considered in previous DNN-based choice models, or used for enhancing behavioral regularity (cf. Wang et al., 2020b; Zheng et al., 2021). In this paper, we address the issue of behavioral irregularity by first defining strong and weak behavioral regularity metrics based on the monotonicity of demand functions, and further designing and implementing a constrained optimization framework that regularizes the input gradient in order to explicitly constrain the gradient’s direction. We then design experiments to examine the performance of behaviorally regularized DNNs in terms of behavioral regularity and predictive performance, differentiating between in-domain and out-of-domain generalizations. We also consider the effect of sample size, which is crucial in practice because large samples are costly for travel surveys. Based on two travel survey datasets from Chicago and London, our experiments compare across the standard DNN and TasteNet (Han et al., 2022) architectures using five evaluation metrics, including log-likelihood, prediction accuracy, $F_{1}$ score, strong behavioral regularity, and weak behavioral regularity. The multinomial logit (MNL) is chosen as a benchmark model due to its concise expression and high behavioral regularity. For TasteNets, we also compare the performance of our gradient regularizers (as soft constraints) and the hard constraints applied in Han et al. (2022). The results show that by using appropriate gradient regularization, both DNNs and TasteNets can achieve high regularity without sacrificing their predictive power, which makes these models competitive in real-world applications and demonstrates the generality of our gradient regularization framework.
@@
The economic choice behavior of humans generally follows the law of demand in economics, which states the inverse relationship between price and quantity demanded. This indicates a monotonic change in market demand due to the change in consumers’ purchasing power, including price and income changes (Chiappori, 1985; Härdle et al., 1991; Hildenbrand, 1983; Quah, 2000). The transportation field has also observed the negative influence of travel costs on travel demand (McFadden, 1974; Souche, 2010; Yao and Morikawa, 2005), based on which demand management policies such as road pricing (May, 1992; Yang and Bell, 1997) were developed. Although such market rationality is widely recognized, individual choice behavior might be irrational (Becker, 1962; Knez et al., 1985). Lichtenstein and Slovic (1971) studied preference reversal in decision making, which is a typical counterexample of individual rationality. Studies in bounded rationality theory (Simon, 1957; Di and Liu, 2016; Watling et al., 2018) and prospect theory (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992) also relax the strict monotonicity assumption in demand modeling.
@@
The law of demand is generally followed by the design of random utility models. In the MNL model, for example, an increase in the travel cost of an alternative would be expected to decrease its systematic utility, thus decreasing its choice probability by design. If the initial model includes estimates with counter-intuitive signs, such as positive cost estimates, an analyst can readily spot them and then refine the model specification or deal with data issues. Once all signs are as expected, the monotonic relationship is guaranteed. By contrast, this is not the case in DNNs because of the complex nonlinear model structure, especially when the number of hidden layers increases. For example, Xia et al. (2023) observed non-monotonic demand predictions with increasing generalized costs in a mode choice experiment with DNNs, which suggests the need to investigate the monotonicity of DNNs to improve their behavioral regularity. Although shallow neural networks (NNs) might reduce the risk of non-monotonic behavior of DNNs (Alwosheel et al., 2019; Zhao et al., 2020), this might come at the cost of reduced modeling flexibility and universal approximation power. Alternatively, Han et al. (2022) and Sifringer et al. (2020) proposed to use DNNs only for learning latent representation in the utility function, while resorting to the DCM framework to ensure model monotonicity. For example, TasteNet, the neural-embedded DCM proposed by Han et al. (2022), assumes a linear model specification with parameters learned by an NN. In this paper, TasteNets are considered as a reference to the standard DNN architecture. On the other hand, depending on model design, these hybrid DNN models might still produce irregular predictions (Wang et al., 2021; Wong and Farooq, 2021). Moreover, hybrid DNN models are a compromise for regularity since they again require the subjective process of model specification. To fully utilize the capability of DNNs, previous studies have attempted to migrate the non-monotonic issue through model ensemble. However, irregular patterns might still be observed after averaging over multiple trainings (Wang et al., 2020a, b; Xia et al., 2023). One promising direction is to directly integrate domain-specific knowledge into the design and training of DNNs, such as incorporating monotonicity constraints in model training (Haj-Yahia et al., 2023). Nevertheless, there is no consensus on how to measure or improve the model regularity of DNNs within the choice modeling field. This paper contributes to the development of a behavioral regularity measure and a novel regularization framework.
@@
As discussed in computer science applications, the regularity of DNNs can be improved by employing either hard or soft constraints. The first category enforces monotonicity by model construction, e.g., constraining the positiveness of weights in hidden layers (Daniels and Velikova, 2010; Dugas et al., 2009; Sill, 1997) through non-negativity constraints (Lawson and Hanson, 1995), restricting the derivatives to be positive (Neumann et al., 2013; Wehenkel and Louppe, 2019), down-weighting samples that violate monotonicity (Archer and Wang, 1993), and incorporating deep lattice networks for learning monotonic functions (You et al., 2017). The second category achieves monotonicity by regularization, i.e., by augmenting a regularization term in the loss function to jointly improve model monotonicity. In the choice modeling domain, recent studies have also incorporated hard constraints on model monotonicity based on the modeler’s prior assumptions to improve intepretability, such as in lattice networks (Kim and Bansal, 2024) and TasteNets (Han et al., 2022). Regularization is firmly rooted in constrained optimization, including the Lagrangian method as an example (Boyd and Vandenberghe, 2004). It has been widely applied as a local method to penalize constraint violations. For example, Sill and Abu-Mostafa (1996) penalized squared deviations in monotonicity for virtual pairs of input variables, while Gupta et al. (2019) proposed a pointwise loss that embeds prior knowledge about monotonicity. Moreover, gradient regularization has also been used to enhance model robustness against adversarial examples (Lyu et al., 2015; Ross and Doshi-Velez, 2018), e.g., penalizing the squared $L_{2}$ norm of the gradient of the loss with respect to (w.r.t.) inputs (Drucker and Le Cun, 1991; Ororbia II et al., 2017), penalizing the squared Frobenius norm of the Jacobian matrix of probabilities (Sokolić et al., 2017) and utilities (Jakubovitz and Giryes, 2018) w.r.t. inputs. Note that regularizing the gradient norms might be less effective to improve monotonicity than regularizing the gradient’s direction like in Haj-Yahia et al. (2023). Inspired by the aforementioned regularization methods, we will design our own approaches in the demand modeling context.
@@
The discrete choice problem is cast as a supervised classification task in DNN-based choice analysis. Assuming there are in total $D$ explanatory variables $\{x_{1},\ldots,x_{D}\}$ for all alternatives, the attribute vector of individual $n$ can be written as $\mathbf{x}_{n}=[x_{n1},\ldots,x_{nD}]^{\top}$. Then, a DNN model predicts the probability of $n$ choosing $i$ out of $J$ alternatives, i.e., $P_{ni}:\mathbb{R}^{D}\to(0,1)$ and $\sum_{i=1}^{J}P_{ni}=1$. The observed choice vector $\mathbf{y}_{n}\in\{0,1\}^{J}$ of $n$ is used for DNN training, where $y_{ni}=1$ if alternative $i$ is chosen, and $y_{ni}=0$ otherwise. Similar to conventional RUMs, DNN models aim to find specifications with high predictive power and behavioral regularity.
In contrast to RUMs with handcrafted specifications, DNNs automatically learn specifications through their unique representation learning capability. Particularly, utility vector $\mathbf{V}_{n}=[V_{n1},\ldots,V_{nJ}]^{\top}$ is specified through a series of transformations, termed as layers $\{f_{1},\ldots,f_{H}\}$, where $H$ denote the total number of layers in a DNN. Each layer $f_{h}$ contains a learnable parameter matrix $W_{h}$, a bias vector $\mathbf{b}_{h}$, and an activation function $\varphi(\cdot)$ (e.g., the rectified linear unit, ReLU) to transform $\mathbf{x}_{n}$. Specifically, each layer transformation can be written as
@@
where $W\in\mathbb{R}^{J\times D}$ can be interpreted as parameters and $\mathbf{b}\in\mathbb{R}^{J\times 1}$ as alternative-specific constants. Although closely related to the MNL model, DNNs allow for flexible model specifications through multi-layer nonlinear transformations. We illustrate in Fig. 1 a feedforward DNN structure with four hidden layers and one classification layer for a choice modeling problem with $D$ attributes for $J$ alternatives.
@@
In this study, we propose a novel metric to evaluate behavioral regularity, which measures the monotonicity of aggregate choice probability functions. The proposed metric essentially measures the monotonicity consistency between the model and prior knowledge on the correct signs of parameter estimates, commonly used in the subjective process of selecting plausible specifications for RUMs. The behavioral regularity metric of alternative $i$ w.r.t. a cost variable $x_{d}$ (with value $x_{nd}$ for individual $n$) is defined as
@@
where $z$ represents a sociodemographic factor of a population group; $P_{i}(\mathbf{x}_{z})$ maps the individual’s attributes $\mathbf{x}_{z}$, including both individual-specific sociodemographics mapped from factor $z$, and alternative-specific cost variables, to the individual’s probability of choosing $i$; $\mathbbm{1}\{\cdot\}$ is an indicator function that equals 1 if ${\partial P_{i}(\mathbf{x}_{z})}/{\partial x_{d}}<\varepsilon$, and 0 otherwise; $\rho(x_{d},z)$ denotes the joint density of attributes $x_{d}$ and the population factor $z$; and $\varepsilon$ represents the modeler’s prior assumptions on the monotonicity of $P_{i}$ w.r.t. $x_{d}$:
@@
  • $\varepsilon>0$, termed as weak regularity, relaxes the strict monotonicity assumption and allows $P_{i}$ to be non-decreasing w.r.t. $x_{d}$. The formulation assumes that some population groups do not respond (with zero derivatives) to $x_{d}$, implying that the behavioral regularity constraint becomes weaker.
As an illustration, a classical linear MNL model with a negative parameter w.r.t. $x_{d}$ yields $B_{id}=1$, which implies that all individual behaviors are consistent with the demand monotonicity assumption and well captured by the model. We also note that Eq. 5 can be interpreted as the cumulative distribution of behavioral regularity with $\varepsilon$ over the whole population and the domain of $x_{d}$.
The population-level behavioral regularity metric in Eq. 5 can be approximated by the mean behavioral regularity across individuals:
@@
where $N$ is the sample size, and the partial derivative is computed with finite differences. The proposed empirical regularity metric in Eq. 6 is the sample analog of the exact metric in Eq. 5. By the Glivenko–Cantelli theorem, Eq. 6 converges in probability to Eq. 5, the cumulative distribution function for the mean population behavioral regularity with $\varepsilon$.
Our behavioral regularity metrics can be extended to incorporate taste heterogeneity across population groups and even individuals within each group by distinguishing $\varepsilon$ w.r.t. different groups, i.e., $\varepsilon$ can be further specified as $\varepsilon_{z}$ to reflect the group-specific thresholds. Meanwhile, our metrics only require aggregate regularity rather than individual regularity, inspired by classical economics discussions: market rationality is a fundamental law, whereas individual behaviors might present more diverse and irrational patterns (Becker, 1962).
@@
The unconstrained formulation in Eq. 7 is sufficient for the estimation of conventional RUMs, since they often satisfy convexity conditions because of linear-in-parameters specifications. In a linear MNL model, for example, choice probability $P_{ni}$ increases monotonically with utility $V_{ni}$ according to Eq. 3. The linear specification in Eq. 4 induces monotonicity of utility w.r.t. cost variables. Therefore, if individuals indeed perceive higher utility with lower costs, the optimization would result in negative parameter estimates, and the behavioral monotonicity is clearly satisfied by evoking the chain rule. The multi-layer nonlinear transformations in DNNs allow for approximation of arbitrary functions, but these complex transformations might lead to non-monotonic choice probability functions, especially when the network is deep. In this case, unconstrained likelihood maximization can no longer guarantee a behaviorally regular model.
@@
where $R$ constrains the attributes $\mathbf{x}_{n}$ of an individual $n$, defined as $R:\mathbb{R}^{J\times D}\to\mathbb{R}$, where dimension $J$ is the number of alternatives and $D$ is the number of attributes. Therefore, the behavioral regularity constraints are imposed at the individual level to enhance the aggregate regularity in $B_{id}$. Specific behavioral regularity constraints will be designed in the next subsection.
@@
Compared with the hard constraint formulation, soft regularization can flexibly accommodate the various degrees of validity in our behavioral regularity assumptions. Similar to the motivation for the weak regularity metric, our approach allows each individual $n$ to somewhat violate the preset constraint $R(\mathbf{x}_{n})$, thus accommodating the potentially irregular behavior of certain individuals. As a result, hyperparameter $\lambda$ provides insight into the consistency between behavioral regularity assumptions and the actual behavior of studied individuals. If a larger $\lambda$ is required to achieve higher predictive performance, it might imply that the actual behavior is inconsistent with prior assumptions, thus providing extra insight into the validity of behavioral regularity constraints. Although the gradient regularization method does not guarantee strong behavioral regularity, it provides insights into model misspecification or the inconsistency between modeling assumptions and data.
@@
For individual $n$, the Jacobian matrix (gradient) of demand vector $\mathbf{P}=[P_{1},\ldots,P_{J}]^{\top}$ w.r.t. explanatory variables $\{x_{1},\ldots,x_{D}\}$ can be written as
@@
  • direct derivatives, e.g., the probability of driving w.r.t. driving cost;
  • cross derivatives, e.g., the probability of driving w.r.t. public transit time; and
  • sociodemographic derivatives, e.g., the probability of driving w.r.t. the age of traveler.